STABILIZED TUNABLE DIODE LASER MEASUREMENTS OF THE P (2) LINE IN THE ^{13}C^{16}O FUNDA-MENTAL BAND BROADENED BY HELIUM AT TEMPERATURES BETWEEN 11.5 AND 298.6 KELVIN

\underline{A. MANTZ}, Department of Physics, Astronomy and Geophysics, Connecticut College, New London, CT 06320; A. HENRY and A. VALENTIN, Laboratoire de Physique Moleculaire et Applications, Universite Pierre et Marie Curie, F-75252 Paris Cedex 05, France.